A merchant for a sporting goods store purchased the following items:

  • 16 dozen Astros baseball hats a cost of $6.00 each
  • 25 dozen Astros baseball t-shirts at a cost of $7.50 each
  • 16 dozen Astros baseball shorts at a cost of $7.75 each

The merchant also received a 5% seasonal discount on the order. The store plans a spring promotion at which all of these Astros items will be sold at the price of $18.99. What is the cumulative markup percentage on this order?

a.

68.93556258%

b.

66.33556258%

c.

64.23556258%

d.

62.43556258%

e.

67.73556258%

f.

None of the above.
Answer #1

Total cost for merchant = 16*6 + 25*7.5 + 16*7.75 = $407.5

5% discount on this so net cost = 407.5*0.95 = $387.125

Total revenue = 18.99(16+25+16) = $1082.43

So, cumulative markup percentage =

So, option C) 64.23556258% is the answer.